OpenStudy (anonymous):

Factor the trinomial: x2 - 10x + 9

OpenStudy (anonymous):

x^2 ----> (x ) (x ) now find what two numbers multiplied and added can equal -10 and +9 -10 + 1 = 9 -10 times + 1 = -10 (x - 10) (x + 1) is your answer

OpenStudy (loser66):

nup

OpenStudy (anonymous):

wat?

OpenStudy (loser66):

it's .... not right